OpenStudy (anonymous):

Find a vector perpendicular to i +2j -k and 3i-j+k. ---> i - 4j -7k (I know how to get this) Hence, find the cartesian equation of the plane parallel to both i +2j-k and 3i-j+k which passes through the point (2,0,-3)

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

(x,y,z)= (2,0,-2)+a(1,2-1)+b(3,-1,1)

OpenStudy (anonymous):

x-2=a+b y = 2a-b z+2 = -a +b eliminate a and b from here and you get the equation

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

How to eliminate a and b?

OpenStudy (anonymous):

adding for example 1º and 3º equation you get: x-2 +z+2 = 2b so b = (x-2 +z +2 )/2 adding 1º and 2º: x-2 +y = 3a so a = (x-2 +y )/3 Now put this values in some equation and put it in the form Ax+By+Cz = D

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

i think if i didn't make any mistakes: x-2y-3z =8 check it just in case...

OpenStudy (anonymous):

ups, i found mistake: x-2=a+b y = 2a-b z+2 = -a +b should be: x-2=a+3b !!!!!!!! y = 2a-b z+2 = -a +b so in the rest fix this value

OpenStudy (anonymous):

result is x-4y-7z =-16 this should be right

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

No.. Not really.... You should get at the end. x-4y-7z=23

OpenStudy (anonymous):

impossible, doesn't contain the point (2,0,-2)

OpenStudy (anonymous):

That's what it says

OpenStudy (anonymous):

try it your self. Put (2,0,-2) in the equation x-4y-7z=23 and it doesn't match it

OpenStudy (anonymous):

Oh, I made a type. It's meant to be (2, 0, -3)

OpenStudy (anonymous):

damm...:)

OpenStudy (anonymous):

ya, than it gives x-4y-7z=23

OpenStudy (anonymous):

@order

OpenStudy (anonymous):

How?

OpenStudy (anonymous):

x-2=a+3b y = 2a-b z+3 = -a +b from here: a =y+z+3, b = (x-2+z+3)/4 puting this values in the last equation: z+3=-y-z-3+ (x-2+z+3)/4 and regrouping gives x-4y-7z=23